### v2

@@ 3.3 Proofs of Theorems 1 and 2
First, we state some lemmas related to the POf and Assumption 1 in Section 3.3.1. Second, we prove Theorem 1 in Section 3.3.2. Third, in Section 3.3.3 we settle the ground for the proof of Theorem 2: this section recalls the convergence analysis of Algorithm 1 from [2, Theorem 1] as Proposition 1 and proves that it is applicable when Assumption 2 holds. Finally, we prove Theorem 2 in Section 3.3.4.
@@
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1. Lemma 1 is related to our notation $\overline{\varphi_{\Omega}}$ and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and Lemma 2 states a basic property of the POf. Lemma 3 allows us, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$). Lemmas 4 and 5 focus on the set $\mathcal{ACC}(\gamma;x)$ involved in Theorems 1.c) and 2. In particular, we prove Theorem 1.c) by using Lemma 5 to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
@@
Let $y\in\mathbb{Y}$ and let $\mathcal{Y}\subseteq\mathbb{Y}$ be a neighbourhood of $y$. Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ be a sequence converging to $y$ and such that $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converges to $z$. Then, for all $k\in\mathbb{N}$ it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$, so we deduce that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$. ∎
@@
[…]
@@
For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$, and $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$. It follows that $\gamma(x_{1})\neq\gamma(x_{2})$. Hence $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection. Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$, and for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)\in\mathbb{Y}(x)$, so $\chi(y)=x$ and then $\gamma(\chi(y))=\gamma(x)=y$. ∎
@@
Consider that Assumption 1 holds. Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$, and $y\in\mathbb{Y}(x)$. Then $\chi(y)=x\in\mathcal{X}$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$. ∎
@@
Under Assumption 1, for all $x\in\mathrm{cl}(X)$ and all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)\subseteq\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ and $\mathcal{ACC}(\gamma;x)$ is finite.
@@
Consider that Assumption 1 holds. Let $x\in\mathrm{cl}(X)$ and let $\mathcal{X}\subseteq\mathbb{X}$ be the neighbourhood of $x$ given by Assumption 1, so $\mathcal{I}\triangleq\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite. For all $i\in\mathcal{I}$, by uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\leavevmode\nobreak\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton, and its unique element denoted by $y_{i}$ satisfies $y_{i}\in\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ since $\gamma(X)\subseteq Y\cap\Omega$ and by continuity of $\chi$. Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}\}$ is a finite subset of $\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$. Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ be a sequence converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$. Then, +$\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}}\}} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$. We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}_{+}\}\subseteq\mathcal{ACC}(\gamma;x)$. ∎
@@
Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. By Lemma 4, it holds that
[…]
so it follows that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ as claimed. ∎
3.3.2 Proof of Theorem 1
Consider that Framework 1 is applicable and that Assumption 1 holds.
@@ 3.3.2 Proof of Theorem 1
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.a).
@@
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(\mathbb{Y}(\mathcal{X}^{*}))\subseteq\mathcal{X}^{*}$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.b).
@@
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so $\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ is nonempty. Select an element $y^{*}$ from it. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 5. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c). ∎
3.3.3 Convergence analysis of Algorithm 1 and proof of validity under Assumption 2

### v3

@@ 3.3 Proofs of Theorems 1 and 2
First, Section 3.3.1 states some lemmas related to the POf and Assumption 1. Second, Section 3.3.2 proves Theorem 1. Third, Section 3.3.3 proves that the convergence analysis of Algorithm 1 from Proposition 1 is applicable when Assumption 2 holds. Finally, Section 3.3.4 proves Theorem 2.
@@ 3.3.1 Technical lemmas on the problem, the framework and its minimal assumption
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{i}}$), the POf, and Assumption 1.
First, Lemma 1 is related to $\overline{\varphi_{\Omega}}$, the extreme barrier variant of $\varphi$ with respect to $\Omega$, and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, the set of accumulation values of $\overline{\varphi_{\Omega}}$ at any point $y\in\mathbb{Y}$. This lemma states that the infimum of $\overline{\varphi_{\Omega}}$ over any neighbourhood of any $y\in\mathbb{Y}$ is below the lowest accumulation value of $\overline{\varphi_{\Omega}}$ at $y$.
@@
Let $y\in\mathbb{Y}$ and let $\mathcal{Y}\subseteq\mathbb{Y}$ be a neighbourhood of $y$. Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. Let $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ converging to $y$ with $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converging to $z$. Then, $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$ holds for all $k\in\mathbb{N}$, so at the limit, $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$ follows. This proves the claim since $z$ was arbitrarily chosen in $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. ∎
Second, Lemma 2 states a basic property of the POf: the oracle function $\gamma$ and the index function $\chi$ induce reciprocal bijections between $X$ and $\gamma(X)$.
@@
In Framework 1, $\gamma$ is a bijection from $X$ to $\gamma(X)$ and its reciprocal is the restriction $\chi_{|\gamma(X)}$.
@@
For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$ by construction of $\gamma$, and $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$ by definition of the partition. It follows that $\gamma(x_{1})\neq\gamma(x_{2})$. Hence, $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection since the surjectivity is clear. Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$ by definition of $\chi$. Finally, for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)$, so $\chi(y)=\chi(\gamma(x))=x$ and then $\gamma(\chi(y))=\gamma(x)=y$. This proves the mutual reciprocity of $\gamma$ and $\chi_{|\gamma(X)}$. ∎
Third, Lemma 3 proves that the union of all partition sets of $\mathbb{Y}$ associated to indices in a neighbourhood of any $x\in\mathbb{X}$ forms a neighbourhood of all $y\in\mathbb{Y}(x)$. We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{i}}$).
@@
Consider that Assumption 1 holds. Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$. For all $y\in\mathbb{Y}(x)$, it holds that $\chi(y)=x\in\mathcal{X}$ by definition of $\chi$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$. This proves the desired claim. ∎
Then, Lemma 4 highlights several key properties resulting from the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1. We prove that these two properties imply that $\mathcal{ACC}(\gamma;x)$ has finitely many elements, which all belong to $\mathbb{Y}(x)$, and that any sequence in $X$ converging to $x$ has its accumulation values through $\gamma$ included in $\mathcal{ACC}(\gamma;x)$. In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
@@
Under Assumption 1, for all $x\in\mathrm{cl}(X)$, the set $\mathcal{ACC}(\gamma;x)$ is a finite subset of $\mathbb{Y}(x)$, and for all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)$.
@@
Consider that Assumption 1 holds. Let $x\in\mathrm{cl}(X)$ and let $\mathcal{X}\subseteq\mathbb{X}$ be the neighbourhood of $x$ given by Assumption 1, so $\mathcal{I}\triangleq\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite. For all $i\in\mathcal{I}$, by completeness of $\mathbb{Y}$ and uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton in $\mathbb{Y}$, and its unique element denoted by $y_{i}$ lies in $\mathbb{Y}(x)$ by continuity of $\chi$. Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\penalty 10000\ i\in\mathcal{I}\}$ is a finite subset of $\mathbb{Y}(x)$. Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$. Then, $\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}\} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$. We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\penalty 10000\ i\in\mathcal{I}_{+}\}\subseteq\{y_{i},\penalty 10000\ i\in\mathcal{I}\}=\mathcal{ACC}(\gamma;x)$, as claimed. ∎
Now, Lemma 5 states a technical equality about how discontinuities of $\gamma$ at any $x\in\mathrm{cl}(X)$ influence the possible accumulation values of $\varphi\circ\gamma$ at $x$. It is an intermediate result to prove Lemma 6 below.
@@
Under Assumption 1, for all $x\in\mathrm{cl}(X)$, it holds that
${\bigcup}_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}}))\quad=\quad{\bigcup}_{y\in\mathcal{ACC}(\gamma;x)}\penalty 10000\ {\bigcup}_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}}).$
Proof.
Consider that Assumption 1 holds, and let $x\in\mathrm{cl}(X)$. We first prove the direct inclusion. Let $z\in\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}})$ and $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$ with $(\varphi(\gamma(x^{k})))_{k\in\mathbb{N}}$ converging to $z$. By Lemma 4, $\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\neq\emptyset$ so there exists $K\subseteq\mathbb{N}$ infinite such that $(\gamma(x^{k}))_{k\in K}$ converges to a limit $y\in\mathcal{ACC}(\gamma;x)$. By re-labelling this sequence by $(y^{k})_{k\in\mathbb{N}}\triangleq(\gamma(x^{k}))_{k\in K}\in\gamma(X)^{\mathbb{N}}$, we observe that $z\in\mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})\subseteq\cup_{y\in\mathcal{ACC}(\gamma;x)}\cup_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})$ as claimed. We now prove the reciprocal inclusion. Let $y\in\mathcal{ACC}(\gamma;x)$ and let $z\in{\bigcup}_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})$. There exists $(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}$ converging to $y$ with $(\varphi(y^{k}))_{k\in\mathbb{N}}$ converging to $z$. By continuity of $\chi$ and Lemma 4, $(x^{k})_{k\in\mathbb{N}}\triangleq(\chi(y^{k}))_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converges to $\chi(y)=x$. Then, $(\varphi(\gamma(x^{k})))_{k\in\mathbb{N}}=(\varphi(y^{k}))_{k\in\mathbb{N}}$ converges to $z$ by Lemma 2, so $z\in{\bigcup}_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}})$ as claimed. ∎
Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) in our proof of Theorem 1.c).
Lemma 6.
Under Assumption 1, $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ holds for all $x\in\mathrm{cl}(X)$.
Proof.
Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. It holds that
@@
It follows that $\inf\mathcal{ACC}(\Phi_{|X};x)\geq\inf_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $\inf\mathcal{ACC}(\Phi;x)=\inf\mathcal{ACC}(\Phi_{|X};x)$ since $\Phi(x^{\prime})=+\infty$ for all $x^{\prime}\notin X$. The proof is completed by finiteness of $\mathcal{ACC}(\gamma;x)$ from Lemma 4, which ensures that $\inf_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)=\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. ∎
3.3.2 Proof of Theorem 1
We now prove Theorem 1. The first two parts consist in a direct verification that given a (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the element $y^{*}\triangleq\gamma(x^{*})$ satisfies the definition of a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{i}}$). This is simple in the global case, while the local case uses Lemma 3 to construct the neighbourhood of $y^{*}$ involved in the definition. Then, the last part where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) works similarly to the (usual) local case, except that it involves one more step to define $y^{*}$. We use Lemma 6 to carefully select $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$, and then we check that $y^{*}$ satisfies the definition of a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
@@ 3.3.2 Proof of Theorem 1
Consider Framework 1 under Assumption 1.
@@
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.a).
@@
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(y)\in\mathcal{X}^{*}$ for all $y\in\mathbb{Y}(\mathcal{X}^{*})$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.b).
@@
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so we may select an element $y^{*}\in\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 6. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.c). ∎
3.3.3 Proof of validity of Proposition 1 under Assumption 2